# Exploring Holomorphic Retracts

### v6

@@ 2.5. Fixed point sets and Retracts
$\partial L\subset\partial B_{1}\times\overline{B}_{2}$. In this case, note that the projection $\pi_{1}:D_{L}\to B_{1}$ is proper and again the general fact that proper holomorphic maps cannot decrease dimension gives ${\rm dim}D_{L}\leq{N_{1}}$. Noting that the arguments in case-(I) leading to (3.3), relied only on the properness of $\pi_{1}$ and not that of $\pi_{2}$, we see that the same arguments apply in this case as well to show again that $D_{L}$ must be of the form
@@
for some $\mathbb{C}$-linear functionals $\beta_{1},\ldots,\beta_{N_{2}}$ on the complex linear subspace $\pi(L)\subset\mathbb{C}^{N_{1}}$; again, the fact that the boundary of $D_{L}$ lies in $B_{1}\times\partial B_{2}$ implies that the linear transformation given by
@@
is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}:=\pi_{1}(D_{L})$ which is the unit ball in $\pi_{1}(L)$ with respect to $\mu_{1}$ isometrically into $B_{2}$; the mapping is surjective if $N_{1}=N_{2}$ and $\pi_{1}(L)=\mathbb{C}^{N}$ where $N=N_{1}=N_{2}$. Picking a point $(w^{0},z^{0})\in\partial D_{L}$, considering the complex line segment joining the origin to this point and the restriction of the retraction of $\rho$ to this segment, we get by the arguments as in Case-(I) by applying the Schwarz lemma of Mazet, that ${\rho^{1}}_{|_{D_{L}}}(w,z)=w$ where $\rho^{1}=\pi_{1}\circ\rho$. As points of $D_{L}$ are of the form $(w,T_{\beta}(w))$ for $w\in\pi_{1}(D_{L})$, we have for all $w\in\pi_{1}(D_{L})$ that
@@
for $w$ varying through $B_{1}^{L}$, the unit ball of $\pi_{1}(L)$ with respect to the norm $\mu_{1}$. Recalling that theorem 3.9 gives $Z=\rho(D_{L})$ whose form we just determined, finishes this case.
@@
$\partial D_{L}\subset\overline{B}_{1}\times\partial B_{2}$. In this case, the projection onto the second-factor namely, $\pi_{2}:D_{L}\to B_{2}$ is proper and we deduce as usual at the outset, that the general fact that proper holomorphic maps do not decrease dimension implies that ${\rm dim}D_{L}\leq N_{2}$.
We then start as in Case-(I) by picking a boundary point $(w^{0},z^{0})\in\partial D_{L}$, which we may assume to not be from the part of $D_{L}$ that lies (if it does at all) in $\partial B_{1}\times\partial B_{2}$ for this has already been dealt with in case-(I); the arguments that follow do not depend on such an assumption and we therefore do not make any assumption about which part of the boundary of $D_{L}$ the point $(w^{0},z^{0})$ comes from. Denote by $l^{\prime}$, the complex line spanned by this point and consider the restriction of our retraction map $\rho$ to the intercept $l^{\prime}\cap D$ of this line with our domain $D=B_{1}\times B_{2}$, which we call $l$. Indeed, compose the standard parametrization of $l$ namely $\varphi(t)=t(w^{0},z^{0})$ with $\rho$ followed by $\pi_{2}$ to get the holomorphic curve in $B_{2}$ given by the parametrization $g_{2}(t):=(\pi_{2}\circ\rho\circ\varphi)(t)$ for $t\in\Delta$. Noting that $g_{2}$ sends $0\in\Delta$ to the origin in $B_{2}$, computing by the chain-rule that $dg_{2}(0)=z^{0}\in\partial B_{2}$ and recalling that all boundary points of $B_{2}$ are $\mathbb{R}$-extremal, we see that all conditions required by theorem 3.4 are met, which when applied to this map $g_{2}$ yields that $g_{2}$ is a linear map: $g_{2}(t)=z^{0}t=tz^{0}$. To see the implications of this, let us write this more elaborately using the original definition of $g_{2}$ as $\pi_{2}(\rho\circ\varphi(t))=tz^{0}$, which unravels explicitly as
@@
meaning that $\rho(D_{L})$ is a $k$-dimensional holomorphic graph over the linear subspace $\pi_{2}(D_{L})$ of $B_{2}$, where $k={\rm dim}(D_{L})={\rm dim}(Z)$. As this verifies in particular that $\rho(D_{L})$ is a submanifold of $Z$, we may now apply the identity principle for analytic sets or just recall theorem 3.9, which allows us to say that the whole of $Z$ is obtained by applying $\rho$ just once to $D_{L}$ i.e., $\rho(D_{L})=Z$; as we just deciphered the form of $\rho(D_{L})$, this finishes determining $Z$. To conclude this case, note that equation (3.6) gives $(\pi_{2}\circ\rho)(0,z)=z$ for all $z\in\pi_{2}(D_{L})$. Moreover, $(\pi_{2}\circ\rho)(D)=\pi_{2}(Z)=\pi_{2}(D_{L})$, since $Z$ is a graph over $\pi_{2}(D_{L})$ i.e., $\pi_{2}\circ\rho$ is a retraction of $D$ with $\pi_{2}(D_{L})$ being the corresponding retract. Restricting $\pi_{2}\circ\rho$ to $B_{2}$ also shows that $\pi_{2}(D_{L})$ is a retract of $B_{2}$ itself. Similar arguments apply to the previous and forthcoming cases as well, which we skip for some conciseness.
@@
This is the final case when $\partial D_{L}$ intersects both $\Gamma_{1}:=\partial B_{1}\times B_{2}$ as well as $\Gamma_{2}:=B_{1}\times\partial B_{2}$. While this does not preclude the possibility that $L$ does not intersect $\Gamma_{0}:=\partial B_{1}\times\partial B_{2}$, certainly any one dimensional complex linear subspace $l^{\prime}$ which intersects $\Gamma_{j}$ for $j=1$ or $2$, intersects the boundary of $D$ only in $\Gamma_{j}$ (note firstly that the $\Gamma_{j}$’s are disjoint). That is to say, if $l=l\cap D$ then $\partial D_{L}$ is fully contained in $\Gamma_{j}$ (for $j=1$ or $2$), as soon as it contains a point from $\Gamma_{j}$. Indeed, if $(w^{0},z^{0})\in\Gamma_{j}$ then the complex line $l$ is the two dimensional real linear subspace spanned by $(w^{0},z^{0})$ and $i\cdot(w^{0},z^{0})=(iw_{0},iz^{0})$. Noting that $w^{0}$ (resp. $z^{0}$) lies in the interior of $\overline{B}_{1}$ (resp. $\overline{B}_{2}$) if and only if $iw^{0}$ (resp. $iz^{0}$) does, verifies the claim that $\partial D_{L}$ is fully contained within $\Gamma_{j}$ (for $j=1$ or $2$) as soon as it intersects $\Gamma_{j}$. As $L$ intersects both $\Gamma_{j}$ for $j=1,2$ we have in particular that $L=T_{0}Z$ cannot equal $l^{\prime}$ and ${\rm dim}(D_{L})\geq 2$. This also means that for each of $\Gamma_{1},\Gamma_{2}$ there are complex lines in $L$ which intersect it. Moreover, if $l$ is a complex line in $D_{L}$ which intersects $\Gamma_{j}$ (for $j=1$ or $2$) then lines obtained by perturbing the coefficients of linear equations which define $L$, slightly, also intersect the open piece $\Gamma_{j}$ (for $j=1$ or $2$) of the boundary of $D$ and by our forgoing remarks, the boundary of the intercept of each such line (near $D_{L}$) with $D$ is fully contained in $\Gamma_{j}$. Observe that the union of these lines contains an open subset $U_{L}$ of $D_{L}$. Next, we apply the Schwarz lemma arguments in Case-(II) (resp. Case-(III)) to each line that intersects $\Gamma_{1}$ (resp. $\Gamma_{2}$). As there are lines in $D_{L}$ which intersect $\Gamma_{1}$ as well as (other) lines in $D_{L}$ which intersect $\Gamma_{2}$, we have both the conclusions of these cases about the major pair of components of $\rho$ namely, $\rho^{j}:=\pi_{j}\circ\rho$ for $j=1,2$, which are that $\rho^{1}$ just returns the first set of variables on $D_{L}$: ${\rho^{1}}_{|_{D_{L}}}(w,z)=w$ and likewise for the second component ${\rho^{2}}_{|_{D_{L}}}(w,z)=z$. Putting together, in short we see that $\rho$ acts just as the identity map, first atleast on $U_{L}$, and thereby on $L$ as well, by the identity principle applied to both the holomorphic maps $\rho^{1},\rho^{2}$. As ${\rm dim}(D_{L})={\rm dim}(Z)$ obviously, this means that $D_{L}=Z$. Therefore, we conclude in this case that $Z$ is a linear subspace of $D$ (of complex dimension $\geq 2$), which is neither contained in $B_{1}$ nor $B_{2}$.
@@ Conclusion
If we drop case-(i), viewing it as a subcase of case (ii) or (iii), then it may be noted that cases (ii) through (iv) are both mutually exclusive and exhaustive as well. Next, it must be noted that each of the possibilities expressed by cases (i) through (iv) are actually realized and can be used to construct the retracts. However, we note that though this theorem seems to provide a complete characterization of the retracts of $D$, we cannot pinpoint which of the linear subspaces of $D$ can happen to be its retract much as we cannot say which linear subspaces of $B_{1}$ or $B_{2}$ are their (linear) retracts as well – this depends on the specific nature of the balls $B_{j}$ i.e., the geometric features of their boundaries, for it is known for certain that not every linear subspace of $D$ can be its retract in general, for instance even in the simplest case $B_{1}=B_{2}=\Delta$, in which case however we can indeed so pinpoint, as was precisely the purpose of our theorem 1.8 above. Finally, we note that the maximum dimension of the nonlinear retracts of $D$ is $\max\{N_{1},N_{2}\}$ far less, in general, than the dimension of $D$ namely, $N_{1}+N_{2}$.
@@
Although $Z$ is biholomorphic to $D_{L_{1}}$, it is unreasonable to expect at-least in general, that there is an automorphism of $D$ which maps $Z$ biholomorphic $D_{L_{1}}$. This fails already for the simplest polyball namely, the polydisc, wherein such a failure can be seen as follows. All the automorphisms of the polydisc are linear fractional transformations, thereby each one of them maps a complex linear subspaces to at-most a quadratic complex surface. Hence, as soon as $Z$ is ‘sufficiently’ non-linear (there are plenty of retracts in the polydisc which are graphs of monomial maps of degree $>2$ for instance, which are herein regarded as ‘sufficiently’ non-linear), it is certainly clear that there does not exist any automorphism of the polydisc mapping $Z$ onto $D_{L_{1}}$. Theorem 1.15 may therefore be viewed as providing an answer to the problem of recasting retracts in a domain into a simple ‘normal’ form (which cannot be reduced to a further simple one in general), by applying some appropriate automorphism of the domain.
@@
The condition of strict convexity is equivalent to the notion of $\mathbb{R}$-extremality (as will be discussed in the next section), which is obtained just by changing the ground field from $\mathbb{C}$ to the reals in that of $\mathbb{C}$-extremality already used in Vesentini’s theorem. The proof of the foregoing theorem will be phrased in a manner to make it clear that as a corollary to its proof it will follow that: we may relax the strict convexity condition on the pair of balls $B_{1},B_{2}$ to $\mathbb{C}$-extremality, to get the same conclusions for those linear retracts $D_{L}$ of $D=B_{1}\times B_{2}$ with the property that some point of $\partial D_{L}$ is an $\mathbb{R}$-extremal point of $\partial D$. This relaxation allows the applicability of this result to the case of balls which are $\mathbb{C}$-extremal but not $\mathbb{R}$-extremal, for which simple examples exist: indeed, the ball in $\mathbb{C}^{N}$ with respect to the $l^{1}$-norm is one such (though balls in all other $l^{p}$-norms for $1<p<\infty$ satisfy the stronger condition of $\mathbb{R}$-extremality.
As mentioned in the introduction, we shall skip the details of the extension of the arguments above to the case of the product of an arbitrary number of balls and only lay down the end result.
@@
Let $B_{1},B_{2},\ldots,B_{N}$ be a collection of bounded balanced strictly convex domains possibly in different complex Euclidean spaces (say, $B_{j}\subset\mathbb{C}^{n_{j}}$ for $j=1,\ldots,N$). Then, every non-trivial retract through the origin of $D=B_{1}\times B_{2}\times\ldots\times B_{N}$ is a holomorphic graph over a linear subspace which itself is a retract of $D$; if the retract is not linear, this linear subspace is infact a subspace of the product of a subcollection of these balls $B_{j}$’s with values in the product of the remaining balls in this collection.
More specifically and precisely, let $L=T_{0}Z$, $D_{L}=L\cap D$ and $\partial D_{L}=L\cap\partial D$. For each $p\in\{1,2,\ldots,N\}$, denote by $\Sigma_{p}$, the set of all $p$-tuples $(j_{1},j_{2},\ldots,j_{p})$ where $\{j_{1},j_{2},\ldots,j_{p}\}\subset\{1,2,\ldots,N\}$ with $j_{1}<j_{2}<\ldots<j_{N}$ (the cardinality of $\Sigma_{p}$ is ($N\choose p$) and note that there exists any given $Z,D_{L}$ as above, $p\in\{1,2,\ldots,N\}$ such that $\partial D_{L}$ is contained in that part of the boundary of the convex body $\overline{D}=\overline{B}_{1}\times\overline{B}_{2}\times\ldots\times\overline{B}_{N}$, obtained by replacing the factor $\overline{B}_{l}$ herein by $\partial B_{l}$ for each $l=j_{1},j_{2},\ldots,j_{p}$ where $\sigma=(j_{1},j_{2},\ldots,j_{p})\in\Sigma_{p}$. Then, $Z$ is the graph of a $B^{\sigma_{c}}$-valued holomorphic map over a $q$-dimensional linear retract of $B_{j_{1}}\times B_{j_{2}}\times\ldots\times B_{j_{p}}$ (which is included in the canonical manner into our product domain $D=B_{1}\times B_{2}\times\ldots\times B_{N}$), where:
@@
    $B^{\sigma_{c}}$ is the product of balls given by $B_{k_{1}}\times B_{k_{2}}\times\ldots\times B_{k_{N-p}}$ where $\sigma_{c}\in\Sigma_{N-p}$ is the $(N-p)$-tuple complementary to $\sigma$ in the sense that $\sigma^{c}=(k_{1},\ldots,k_{N-p})\subset\{1,\ldots,N\}\setminus\{j_{1},\ldots,j_{p}\}$ and $k_{1}<k_{2}<\ldots<k_{N-p}$.
@@
The simplest case for $D$ as in the above theorem is obtained by taking all the $B_{j}$’s therein to be just the unit disc, in which case ofcourse $D$ is the standard unit polydisc $\Delta^{N}$, whose retracts were determined originally by Heath and Suffridge [Heath_Suff]. To recover their result from the above theorem which only characterizes the retracts through the origin, we only have to note that the automorphism group of $\Delta^{N}$ acts transitively thereby implying that all retracts of $\Delta^{N}$ are obtained as biholomorphic images of those passing through the origin under some automorphism of $\Delta^{N}$; recalling that all automorphisms of $\Delta^{N}$ are actually fractional linear maps, we also note that the main essential feature of retracts through the origin described in the above theorem namely, as holomorphic graphs over some linear subspace is inherited by all retracts of $\Delta^{N}$, as fractional linear maps carry linear subspaces to affine-linear subspaces.
Similar considerations apply whenever the factors $B_{j}$ are homogeneous domains; recall that a domain is said to be homogeneous if its automorphism group acts transitively and examples of bounded balanced convex domains which are also homogeneous include apart from $\Delta^{N}$, the standard Euclidean ball $\mathbb{B}^{N}$ and more generally, products of Cartan domains of various types (better known as bounded balanced symmetric domains). In such a case their product $D=B_{1}\times\ldots B_{N}$ is homogeneous as well and all its retracts are obtained as images under some element of ${\rm Aut}(D)$, of those retracts of $D$ passing through the origin which have been determined by the above theorem. However, we may not be able to say that every such retract (unless it passes through the origin), is a holomorphic graph over a linear subspace since elements of ${\rm Aut}(D)$ may well not be so simple (as linear fractional transformations) to preserve linearity as in the aforementioned instance of the polydisc. To give another concrete instance, where the main features of the retracts through the origin are indeed preserved, let us take each $B_{j}$ in the above theorem to be the standard unit ball with respect to the $l^{2}$-norm; so, let $\mathbb{B}_{1},\ldots,\mathbb{B}_{N}$ be standard Euclidean balls in complex Euclidean spaces of possibly different dimensions. Then as ${\rm Aut}(D)={\rm Aut}(\mathbb{B}_{1})\times\ldots\times{\rm Aut}(\mathbb{B}_{N})$, we see that the product domain $D=\mathbb{B}_{1}\times\ldots\times\mathbb{B}_{N}$ is also homogeneous and more importantly, all of its automorphisms are fractional-linear maps. However, despite all retracts of each of these $l^{2}$-balls being affine-linear, this does not mean that every retract of $D$ is affine-linear because it must be kept in mind that the operation of taking Cartesian product leads to new retracts in the form of holomorphic graphs (even through the origin), which are not linear.
The foregoing remarks for a product of Euclidean balls apply as well to the case of products of balls with polydiscs such as $\Delta^{k}\times\mathbb{B}$ (or a larger collection of Euclidean balls and polydiscs in any order) which are also homogeneous domains with all of is automorphisms being fractional-linear maps. We conclude by noting that in these cases the characterization of retracts provided by theorems 1.15 and 3.20, is complete.
@@
The proof follows ideas from the artice [Fost] of Fornaess–Sibony. Let us begin by noting that it follows from theorem 1.13 proved in the foregoing section that every retract of a bounded balanced homogeneous domain is biholomorphic to the graph of some holomorphic function over a linear subspace of $D$.

### v7

@@ 2.9. Analytic and Semi-analytic sets
$\partial D_{L}\subset\partial B_{1}\times\overline{B}_{2}$. In this case, note that the projection $\pi_{1}:D_{L}\to B_{1}$ is proper and again the general fact that proper holomorphic maps cannot decrease dimension gives ${\rm dim}D_{L}\leq{N_{1}}$. Noting that the arguments in case-(I) leading to (3.3), relied only on the properness of $\pi_{1}$ and not that of $\pi_{2}$, we see that the same arguments apply in this case as well to show again that $D_{L}$ must be of the form
@@
for some $\mathbb{C}$-linear functionals $\beta_{1},\ldots,\beta_{N_{2}}$ on the complex linear subspace $\pi_{1}(L)\subset\mathbb{C}^{N_{1}}$; again, the fact that the boundary of $D_{L}$ lies in $\partial B_{1}\times B_{2}$ implies that the linear transformation given by
@@
maps $B_{1}^{L}:=\pi_{1}(D_{L})$ into $B_{2}$; this may be phrased in the convex case as: $T$ is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into ($\mathbb{C}^{N_{2}}$, $\mu_{2}$). The mapping is surjective if $N_{1}=N_{2}$ and $\pi_{1}(L)=\mathbb{C}^{N}$ where $N=N_{1}=N_{2}$. Picking a point $(w^{0},z^{0})\in\partial D_{L}$, considering the complex line segment joining the origin to this point and the restriction of the retraction of $\rho$ to this segment, we get by the arguments as in Case-(I) by applying theorem 3.6 (or in the convex case, the Schwarz lemma of Mazet), that ${\rho^{1}}_{|_{D_{L}}}(w,z)=w$ where $\rho^{1}=\pi_{1}\circ\rho$. As points of $D_{L}$ are of the form $(w,T_{\beta}(w))$ for $w\in\pi_{1}(D_{L})$, we have for all $w\in\pi_{1}(D_{L})$ that
@@
for $w$ varying through $B_{1}^{L}$, the unit ‘ball’ of $\pi_{1}(L)$ with respect to the functional $\mu_{1}$. Recalling that theorem 3.10 gives $Z=\rho(D_{L})$ whose form we just determined, finishes this case, yielding that $Z$ is a graph of a $B_{2}$-valued holomorphic map on the linear subspace $B_{1}^{L}$, which we also denote by $(B_{1})_{L_{1}}$.
@@
$\partial D_{L}\subset\overline{B}_{1}\times\partial B_{2}$. In this case, the projection onto the second-factor namely, $\pi_{2}:D_{L}\to B_{2}$ is proper and we deduce as usual at the outset, that: the general fact that proper holomorphic maps do not decrease dimension implies that ${\rm dim}D_{L}\leq N_{2}$.
We then start as in Case-(I) by picking a boundary point $(w^{0},z^{0})\in\partial D_{L}$, which we may assume to not be from the part of $D_{L}$ that lies (if it does at all) in $\partial B_{1}\times\partial B_{2}$ for this has already been dealt with in case-(I); the arguments that follow do not depend on such an assumption and we therefore do not make any assumption about which part of the boundary of $D_{L}$ the point $(w^{0},z^{0})$ comes from. Denote by $l^{\prime}$, the complex line spanned by this point and consider the restriction of our retraction map $\rho$ to the intercept $l^{\prime}\cap D$ of this line with our domain $D=B_{1}\times B_{2}$, which we call $l$. Indeed, compose the standard parametrization of $l$ namely $\varphi(t)=t(w^{0},z^{0})$ with $\rho$ followed by $\pi_{2}$ to get the holomorphic curve in $B_{2}$ given by the parametrization $g_{2}(t):=(\pi_{2}\circ\rho\circ\varphi)(t)$ for $t\in\Delta$. Noting that $g_{2}$ sends $0\in\Delta$ to the origin in $B_{2}$, computing by the chain-rule that $dg_{2}(0)=z^{0}\in\partial B_{2}$ and recalling that all boundary points of $B_{2}$ are holomorphically extreme, we see that all conditions required by theorem 3.6 (or if one so wishes, the Mazet Schwarz lemma for the convex case) are met; which when applied to $g_{2}$ yields that this map is a linear map: $g_{2}(t)=z^{0}t=tz^{0}$. To see the implications of this, let us write this more elaborately using the original definition of $g_{2}$ as $\pi_{2}(\rho\circ\varphi(t))=tz^{0}$, which unravels explicitly as
@@
meaning that $\rho(D_{L})$ is a $k$-dimensional holomorphic graph over the linear subspace $\pi_{2}(D_{L})$ of $B_{2}$, where $k={\rm dim}(D_{L})={\rm dim}(Z)$. As this verifies in particular that $\rho(D_{L})$ is a submanifold of $Z$, we may now apply the identity principle for analytic sets or just recall theorem 3.10, which allows us to say that the whole of $Z$ is obtained by applying $\rho$ just once to $D_{L}$ i.e., $\rho(D_{L})=Z$; as we just deciphered the form of $\rho(D_{L})$, this finishes determining $Z$.
@@
This is the final case when $\partial D_{L}$ intersects both $\Gamma_{1}:=\partial B_{1}\times B_{2}$ as well as $\Gamma_{2}:=B_{1}\times\partial B_{2}$. While this does not preclude the possibility that $L$ does not intersect $\Gamma_{0}:=\partial B_{1}\times\partial B_{2}$, certainly any one dimensional complex linear subspace $l^{\prime}$ which intersects $\Gamma_{j}$ for $j=1$ or $2$, intersects the boundary of $D$ only in $\Gamma_{j}$ (note firstly that the $\Gamma_{j}$’s are disjoint). That is to say, if $l=l\cap D$ then $\partial D_{L}$ is fully contained in $\Gamma_{j}$ (for $j=1$ or $2$), as soon as it contains a point from $\Gamma_{j}$. Indeed, if $(w^{0},z^{0})\in\Gamma_{j}$ then the complex line $l$ is the two dimensional real linear subspace spanned by $(w^{0},z^{0})$ and $i\cdot(w^{0},z^{0})=(iw_{0},iz^{0})$. Noting that $w^{0}$ (resp. $z^{0}$) lies in the interior of $\overline{B}_{1}$ (resp. $\overline{B}_{2}$) if and only if $iw^{0}$ (resp. $iz^{0}$) does, verifies the claim that $\partial D_{L}$ is fully contained within $\Gamma_{j}$ (for $j=1$ or $2$) as soon as it intersects $\Gamma_{j}$. As $L$ intersects both $\Gamma_{j}$ for $j=1,2$ we have in particular that $L=T_{0}Z$ cannot equal $l^{\prime}$ and ${\rm dim}(D_{L})\geq 2$. This also means that for each of $\Gamma_{1},\Gamma_{2}$ there are complex lines in $L$ which intersect it. Moreover, if $l$ is a complex line in $D_{L}$ which intersects $\Gamma_{j}$ (for $j=1$ or $2$) then lines obtained by perturbing the coefficients of linear equations which define $L$, slightly, also intersect the open piece $\Gamma_{j}$ (for $j=1$ or $2$) of the boundary of $D$ and by our forgoing remarks, the boundary of the intercept of each such line (near $D_{L}$) with $D$ is fully contained in $\Gamma_{j}$. Observe that the union of these lines contains an open subset $U_{L}$ of $D_{L}$. Next, we apply the Schwarz lemma arguments in Case-(II) (resp. Case-(III)) to each line that intersects $\Gamma_{1}$ (resp. $\Gamma_{2}$). As there are lines in $D_{L}$ which intersect $\Gamma_{1}$ as well as (other) lines in $D_{L}$ which intersect $\Gamma_{2}$, we have both the conclusions of these cases about the major pair of components of $\rho$ namely, $\rho^{j}:=\pi_{j}\circ\rho$ for $j=1,2$, which are that $\rho^{1}$ just returns the first set of variables on $D_{L}$: ${\rho^{1}}_{|_{D_{L}}}(w,z)=w$ and likewise for the second component ${\rho^{2}}_{|_{D_{L}}}(w,z)=z$. Putting together, we therefore see that $\rho$ acts just as the identity map, first atleast on $U_{L}$, and thereby on $L$ as well, by the identity principle applied to both the holomorphic maps $\rho^{1},\rho^{2}$. As ${\rm dim}(D_{L})={\rm dim}(Z)$ obviously, this means that $D_{L}=Z$. Therefore, we conclude in this case that $Z$ is a linear subspace of $D$ (of complex dimension $\geq 2$), which is neither contained in $B_{1}$ nor $B_{2}$.
@@ Conclusion
If we drop case-(i), viewing it as a subcase of case (ii) or (iii), then it may be noted that cases (ii) through (iv) are both mutually exclusive and exhaustive as well. Next, it must be noted that each of the possibilities expressed by cases (i) through (iv) are actually realized and can be used to construct the retracts. However, we note that though this theorem seems to provide a complete characterization of the retracts of $D$, we cannot pinpoint which of the linear subspaces of $D$ can happen to be its retract much as we cannot say which linear subspaces of $B_{1}$ or $B_{2}$ are their (linear) retracts as well – this depends on the specific nature of the balls $B_{j}$ i.e., the geometric features of their boundaries, for it is known for certain that not every linear subspace of $D$ can be its retract in general, for instance even in the simplest case $B_{1}=B_{2}=\Delta$, in which case however we can indeed so pinpoint, as was precisely the purpose of our theorem 1.7 above.
@@
Although $Z$ is biholomorphic to $D_{L_{1}}$, it is unreasonable to expect at-least in general, that there is an automorphism of $D$ which maps $Z$ biholomorphic $D_{L_{1}}$. This fails already for the simplest polyball namely, the polydisc, wherein such a failure can be seen as follows. All the automorphisms of the polydisc are linear fractional transformations, thereby each one of them maps a complex linear subspaces to at-most a quadratic complex surface. Hence, as soon as $Z$ is ‘sufficiently’ non-linear (there are plenty of retracts in the polydisc which are graphs of monomial maps of degree >2>2 for instance, which are herein regarded as ‘sufficiently’ non-linear), it is certainly clear that there does not exist any automorphism of the polydisc mapping $Z$ onto $D_{L_{1}}$. Theorem 1.17 may therefore be viewed as providing an answer to the problem of recasting retracts in a domain into a simple ‘normal’ form (which cannot be reduced to a further simple one in general), by applying some appropriate automorphism of the domain.
@@
It may very well happen that neither $\pi_{1}(Z),\pi_{2}(Z)$ need be linear subspaces of the domain, let alone retracts (neither of the respective factors $B_{1},B_{2}$ nor of $D$). For a simple example, consider the polydisc in $\mathbb{C}^{4}$ which we view as a product of $2$ copies of the bidisc i.e., let $D:=\Delta^{2}\times\Delta^{2}$; and, consider the linear subspace $L$ of $\mathbb{C}^{4}$ defined by $L:=span\{v_{1},v_{2}\}$, where $v_{1}=(1,1/2,1/4,0)\in\partial\Delta^{2}\times\Delta^{2},~v_{2}=(0,1/2,3/4,1)\in\Delta^{2}\times\partial\Delta^{2}$. We can identify first and second factor of $D$ as $\Delta^{2}\times\{0\}$ and $\{0\}\times\Delta^{2}$. Note that the holomorphic map $\rho$ on $D$ defined by $\rho(z_{1},z_{2},z_{3},z_{4})=(z_{1},\frac{z_{1}+z_{4}}{2},\frac{z_{1}+3z_{4}}{4},z_{4})$ gives a retraction map from $D$ onto $D_{L}$. Note that $D_{L}$ is a retract of $D$ and $\partial D_{L}$ intersects $\partial\Delta^{2}\times\Delta^{2}$ and $\Delta^{2}\times\partial\Delta^{2}$. Now we consider the projections restricted to $D_{L}$, which is given as follows:
$\pi_{1}:L\to\pi_{1}(L)$ defined by
$\pi_{1}(z_{1},\frac{z_{1}+z_{4}}{2},\frac{z_{1}+3z_{4}}{4},z_{4})=(z_{1},\frac{z_{1}+z_{4}}{2},0,0).$
and $\pi_{2}:L\to\pi_{2}(L)$ defined by
$\pi_{2}(z_{1},\frac{z_{1}+z_{4}}{2},\frac{z_{1}+3z_{4}}{4},z_{4})=(0,0,\frac{z_{1}+3z_{4}}{4},z_{4}).$
Note that $\pi_{1}(D_{L})$ and $\pi_{2}(D_{L})$ are not linear subspaces of neither the respective factors ($\Delta^{2}\times\{0\}$, $\{0\}\times\Delta^{2}$) nor of $D$; indeed, this can be confirmed for instance by noting that that $(0,1/2)=\pi_{1}(v_{2})\in\pi_{1}(D_{L})$ and $(1/4,0)=\pi_{2}(v_{1})\in\pi_{2}(D_{L})$, but their scalings: $(0,3/4)\notin\pi_{1}(D_{L})$ and $(1/2,0)\notin\pi_{2}(D_{L})$ despite $(0,3/4,0,0)$ and $(1/2,0,0,0)$ are within $D$.
As mentioned in the introduction, we shall skip the details of the extension of the arguments above to the case of an extended product of an arbitrary number of ‘balls’ and only lay down the end result; in-fact we only do this for the case of convex factors with $\mathbb{C}$-extremal boundaries which is of special importance as it leads to an associated result about norm-one projections with respect to those norms whose unit balls are such products. The corresponding statement for bounded balanced pseudoconvex domains with holomorphically extreme boundaries is also true and can be written down as in theorem 1.17.
@@
Let $B_{1},B_{2},\ldots,B_{N}$ be a collection of bounded balanced convex domains with $\mathbb{C}$-extremal boundaries, possibly in different complex Euclidean spaces (say, $B_{j}\subset\mathbb{C}^{n_{j}}$ for $j=1,\ldots,N$). Then, every non-trivial retract through the origin of $D=B_{1}\times B_{2}\times\ldots\times B_{N}$ is a holomorphic graph over a linear subspace which itself is a retract of $D$; if the retract is not linear, this linear subspace is infact a subspace of the product of a subcollection of these balls $B_{j}$’s with values in the product of the remaining balls in this collection.
[…]
@@
    $B^{\sigma_{c}}$ is the product of balls given by $B_{k_{1}}\times B_{k_{2}}\times\ldots\times B_{k_{N-p}}$ where $\sigma_{c}\in\Sigma_{N-p}$ is the (Np)(N-p)-tuple complementary to $\sigma$ in the sense that $\sigma^{c}=(k_{1},\ldots,k_{N-p})\subset\{1,\ldots,N\}\setminus\{j_{1},\ldots,j_{p}\}$ and $k_{1}<k_{2}<\ldots<k_{N-p}$.
@@
Needless to say, when the product domain happens to be homogeneous such as a bounded balanced symmetric domain whose automorphisms are known, the above theorem suffices to give all its retracts.
@@
The proof follows ideas from the article [Fost] of Fornæss–Sibony. Let us begin by noting that it follows from theorem 1.15 proved in the foregoing section that every retract of a bounded balanced homogeneous domain is biholomorphic to the graph of some holomorphic function over a linear subspace of $D$.